# Multimodal Transformer Distillation for Audio-Visual Synchronization

### 20 Dec 2023 v2

@@ 2.1 VocaLiST
VocaLiST [13] is a SOTA audio-visual synchronization model. The input of VocaLiST is a sequence of visuals and its corresponding audio features. The output is about to classify whether a given audio-visual pair is in sync. VocaLiST consists of an audio-visual front-end and a synchronization back-end. The audio-visual front end extracts audio and visual features. The synchronization back-end comprises three cross-modal Transformer encoders, namely audio-to-visual (AV) Transformer, visual-to-audio (VA) Transformer, and Fusion Transformer. Each Transformer block has 4 layers. The core part of the cross-modal Transformer is the cross-attention layer, whose input has queries, keys, and values. In VocaLiST, the AV Transformer uses audio features for queries and visual features for keys and values, while the VA Transformer does the opposite. The Fusion Transformer combines these, using the AV output for its queries and the VA output for keys and values. Its output is max-pooled over time dimension and activated via tanh function. Finally, a fully connected layer serves as a classifier, determining if voice and lip motion are synchronized, with optimization through binary cross-entropy loss.

### 18 Mar 2024 v3

@@ 2.1 VocaLiST
VocaLiST [13] is a SOTA audio-visual synchronization model. The input of VocaLiST is a sequence of visuals and its corresponding audio features. The output is about to classify whether a given audio-visual pair is in sync. VocaLiST consists of an audio-visual front-end and a synchronization back-end. The audio-visual front end extracts audio and visual features. The synchronization back-end comprises three cross-modal Transformer encoders, namely audio-to-visual (AV) Transformer, visual-to-audio (VA) Transformer, and Fusion Transformer. Each Transformer block has 4 layers. The core part of the cross-modal Transformer is the cross-attention layer, whose input has queries, keys, and values. In VocaLiST, the AV Transformer uses audio for queries and visual data for keys and values, while the VA Transformer does the opposite. The Fusion Transformer merges these, taking AV output for queries and VA output for keys and values. Its output undergoes max-pooling over time and is activated with a tanh function. A fully connected layer then classifies if voice and lip motion are synchronized, using binary cross-entropy loss for optimization.